nail polish remover is a mix of water & acetone. How much acetone must be combined with a solution that is 28% acetone to make 24 oz of a 58% solution?

Answer by checkley71(8403)

.28X+(24-X)=.58*24
.28X+24-X=13.92
-.72X=13.92-24
.72X=10.08
X=10.08/.72
X=14 OZ. OF 28% ACETONE IS REQUIRED.
THEREFORE 24-14=10 OZ. OF PURE ACETONE IS ALSO REQUIRED.
PROOF:
.28*14+24-14=13.92
3.92+10=13.92
13.92=13.92